The auto repair shop of Quality Motor Company uses standards to control the labor time and labor cost in the shop. The standard labor cost for a motor tune-up is given below: Job Standard Hours: 3.00, Standard Rate: $7.00, Standard Cost: $21.00. The record showing the time spent in the shop last week on motor tune-ups has been misplaced. However, the shop supervisor recalls that 250 tune-ups were completed during the week, and the controller recalls the following variance data relating to tune-ups:

Labor rate variance $ 1,020 F Labor spending variance $ 2,700 F

Required: 1. Determine the number of actual labor-hours spent on tune-ups during the week.
Answer #1
Labour spending Variance = Labour Rate Variance+Labour Efficiency Variance
Labor Efficiency Variance = Labor Spending Variance -Labor Rate Variance
Labor Efficiency Variance = 2,700U - $1020F
=1,680 F
Requirement 1
Computation of Actual labor hours spent on tune-ups
Labor Efficiency Variance = (Standard Hours for Actual Output - Actual Hours worked ) X Standard rate
1680 F = ((250x3) -AH) X 7.00
1680 /7.00 = 750-AH
AH =750-240
AH =510 hours